### v1

# WWW MIR (2025)

[…]
WWW MIR (2025)
A Multi-Granularity Multimodal Retrieval Framework for Multimodal Document Tasks
Mingjun Xu1    Zehui Wang1    Hengxing Cai1,2    Renxin Zhong2
( 1 DP Technology
2 Sun Yat-sen University )
1 Task Description
Retrieval-augmented generation (RAG) systems are widely adopted to enhance the factual consistency of large language models. However, most existing RAG systems are limited to text-only retrieval, making them ineffective for visually-rich documents that contain a combination of text, images, tables, and charts. To bridge this gap, the Multimodal Document Retrieval (MDR) Challenge [^1] https://erel-mir.github.io/challenge/mdr-track1/ introduces two benchmark tasks:
  • Task 1: MMDocIR – Multimodal Page Retrieval, which focuses on retrieving full pages from long documents across various domains such as industry reports and academic papers.
  • Task 2: M2KR – Open-Domain Visual Retrieval, which involves identifying relevant image regions or pages from visually-rich web content like Wikipedia.
Both tasks require a unified retrieval system capable of modeling complex interdependencies across textual and visual modalities, handling heterogeneous document layouts, and maintaining fine-grained relevance under long-context settings. Performance is measured using average recall at top-1, top-3, and top-5 rankings [fu20251sterelmirworkshopefficient].
Our solution, developed by Team LLMHunter, adopts a training-free hybrid framework that integrates multiple off-the-shelf text and vision-language models in a modular design, enabling robust retrieval without task-specific fine-tuning. By combining diverse retrieval signals—such as semantic text similarity, OCR-enhanced image-text alignment, and layout-aware heuristics—our method emphasizes robustness, zero-shot generalization, and ease of deployment. Without relying on task-specific fine-tuning, our approach achieves competitive performance and establishes a strong baseline for multimodal document retrieval.
We detail the overall architecture, model choices, fusion strategies, and experimental insights that guided our design, and discuss how training-free multimodal retrieval can provide a scalable and reproducible solution to real-world multimodal information retrieval challenges.
2 Related Work
Multimodal document retrieval has garnered significant attention, leading to the development of models that effectively integrate visual and textual information. ColPali [faysse2025colpaliefficientdocumentretrieval] introduces a vision-language model that directly embeds document page images into high-quality multi-vector representations, utilizing a late interaction mechanism for efficient retrieval. This approach simplifies the retrieval pipeline by eliminating the need for text extraction and layout analysis, achieving superior performance on the ViDoRe benchmark. Complementing this, GME [zhang2025gmeimprovinguniversalmultimodal] presents a universal multimodal retrieval framework leveraging multimodal large language models (MLLMs). By constructing a large-scale, high-quality fused-modal training dataset, GME addresses modality imbalance and enables the embedding of diverse inputs—text, images, and their combinations—into a unified vector space, demonstrating state-of-the-art performance on the Universal Multimodal Retrieval Benchmark (UMRB).
To evaluate and benchmark such retrieval systems, MMDocIR [dong2025mmdocirbenchmarkingmultimodalretrieval] offers a comprehensive dataset comprising long documents with annotations for both page-level and layout-level retrieval tasks. This benchmark facilitates the assessment of systems’ abilities to retrieve relevant content across various modalities and granularities. Similarly, M2KR [lin2024preflmrscalingfinegrainedlateinteraction] provides a suite of datasets designed for training and evaluating general-purpose vision-language retrievers across multiple tasks, including image-to-text and question-to-text retrieval. M2KR serves as a valuable resource for developing and benchmarking models capable of handling diverse multimodal retrieval scenarios.
3 Methods
Our overall approach follows a multi-stage retrieval framework based on multi-granularity modeling. Given a text or multimodal query, we first encode both queries and document candidates into a unified embedding space across different levels of granularity, including page-level and region-level representations. Retrieval is initially conducted through similarity-based matching, and fusion strategies are employed to integrate multiple retrieval signals. Finally, reranking modules are used to refine the candidate results for enhanced precision.
Both the M2KR and MMDocIR tasks utilize this multi-granularity design philosophy, with M2KR focusing on fine-grained region retrieval and MMDocIR emphasizing full-page retrieval combined with hierarchical reranking. In the following subsections, we present detailed descriptions of our methods for each track.
3.1 Overall Retrieval Framework
To address the challenges of multimodal document retrieval, we design a comprehensive retrieval framework that unifies multi-granularity representations and multimodal fusion strategies. The framework is structured into two main components: (1) M2KR, focusing on fine-grained region-level retrieval through integrated visual and textual signals; and (2) MMDocIR, which combines full-page retrieval with detailed regional and OCR-based retrieval stages. Both components share a core architecture of multi-stage matching, similarity-based scoring, and reranking modules to ensure accurate and robust retrieval across diverse document types and query formats. The following sections provide detailed descriptions of these two task-specific solutions.
3.1.1 M2KR: Multimodal Region Retrieval with Fusion
In the M2KR task, we propose a fine-grained multimodal retrieval framework that integrates visual and textual signals at the region level. By leveraging multi-granularity representations, our method effectively captures the detailed layout and semantic information within document pages to enhance retrieval performance.
Figure 1: Overview of the Multi-Granularity Multimodal Retrieval Framework for M2KR.
Input: In the M2KR task, the retrieval process starts by preparing multi-granularity inputs. As is shown in the Figure 1 . For each candidate page, we segment it into multiple regions to obtain region-level representations. Simultaneously, text descriptions for the candidates are generated using the Qwen2.5-VL model, providing complementary semantic information. The user query, which can be in text, image, or multimodal format, is also processed into a unified embedding space.
Process: During retrieval, three types of matching strategies are applied: pure image retrieval based on region embeddings, multimodal retrieval combining query and region features, and pure text retrieval between query and candidate text descriptions. Cosine similarity (CosSim) is calculated for each modality. The resulting relevance scores from the three matching strategies are then fused to jointly consider visual, multimodal, and textual signals.
Output: After score fusion, a reranking module is employed to refine the candidate results. The fusion strategy effectively balances different modalities, while the final rerank step ensures the retrieval results are aligned with the true query intent.
3.1.2 MMDocIR: Full-Page Retrieval with Multistage Ranking
In the MMDocIR task, we design a hierarchical retrieval framework that combines full-page and regional information. By integrating global page-level retrieval with fine-grained regional matching and OCR text retrieval, our method achieves robust document understanding and significantly improves retrieval accuracy.
Figure 2: Overview of the MMDocIR retrieval framework combining full-page and regional information.
Input: In the MMDocIR task, the retrieval process begins with a user query provided in text form. As is shown in the Figure 2. Candidate document pages are prepared with their full-page images, segmented region images, and OCR-recognized texts to enable multi-granularity retrieval.
Process: First, ColQwen2-7B is used to encode the user query and the full-page image candidates to perform initial full-page retrieval. Simultaneously, we apply GME-7B to conduct two additional retrieval stages: (1) text-to-OCR-text retrieval, where the query is matched against the OCR texts extracted from the pages; and (2) text-to-region-image retrieval, where the query is matched against segmented regional images. Cosine similarity (CosSim) is calculated for each retrieval path. The relevance scores obtained from full-page retrieval, OCR-text retrieval, and regional-image retrieval are then fused to combine global, textual, and regional information.
Output: After score fusion, a Qwen2.5-VL Reranker is employed to perform fine-grained reranking over the fused candidate results. This reranking module further refines retrieval accuracy by leveraging cross-modal understanding, producing the final retrieval output.
3.2 Fusion and Reranking Mechanism
To maximize retrieval accuracy, we employ a combined strategy of multi-source score fusion and reranking. This ensures that both diverse retrieval paths and final semantic alignment are fully leveraged.
3.2.1 Multi-Source Score Fusion
We implement a two-layer score fusion process. First, in the modality score integration stage, the GME model generates three types of relevance scores for each candidate: text-to-OCR-text similarity, text-to-region-image similarity, and multimodal similarity between the query and region features. These scores capture different aspects of document relevance and are directly used for initial ranking, providing a rich multi-granularity matching signal.
Second, to combine the outputs from both the GME and ColQwen retrieval paths, we apply Reciprocal Rank Fusion (RRF). This method merges ranked lists by assigning higher weight to top-ranked results from each retrieval path, effectively balancing global and fine-grained retrieval signals. The fused ranking ensures that strong candidates from different models are jointly considered, enhancing retrieval robustness.
3.2.2 Reranking with Qwen-VL
After the fusion step, we apply a reranking process using the Qwen2.5-VL model. This reranker performs a binary relevance judgment, predicting whether each candidate is a correct match (labeled as Yes) or not (labeled as No). Candidates classified as Yes are then promoted to the top of the final ranking list. This reranking step ensures that high-confidence matches are prioritized, further refining retrieval accuracy through cross-modal semantic alignment.
4 Results and Analysis
In this section, we present the experimental results of different retrieval configurations on the M2KR and MMDocIR tasks. We evaluate the progressive improvements brought by integrating advanced modules such as ColQwen2.5, layout-aware search, and reranking strategies. Table 1 and Figure 3 summarize the performance of each variant in terms of retrieval score.
Table 1: Ablation results on M2KR and MMDocIR tasks.
Method Score
GME Baseline 53.9613
GME for M2KR, ColQwen2.5 for MMDocIR 56.6967
GME Instruct Search + ColQwen2-7B (bugfix) 60.3278
MMDocIR Rerank 61.5289
GME for M2KR Layout Search 64.0902
M2KR Rerank 65.5588
Figure 3: Performance trend across progressive improvements.
The results show a clear upward trend as we incorporate task-specific enhancements. Applying ColQwen2.5 and fixing multimodal bugs led to a substantial boost in MMDocIR, while introducing layout-aware search and reranking significantly benefited M2KR. Our final configuration, which integrates reranking for both tasks, achieves the highest retrieval score of 65.56.
5 Conclusion
In this work, we propose a unified multi-granularity multimodal retrieval framework, and apply it to two document-level tasks: M2KR and MMDocIR. By leveraging hierarchical encoding, modality-aware retrieval strategies, and reranking mechanisms, our approach progressively enhances retrieval accuracy. Experimental results show a clear performance gain from the integration of layout-aware search and reranking modules, with the final system achieving a score of 65.56. This demonstrates the effectiveness and scalability of our framework in complex document retrieval scenarios.
[…]